# SPFFNet: Strip Perception and Feature Fusion Spatial Pyramid Pooling for Fabric Defect Detection

### v1

@@ Abstract
Defect detection in fabrics is critical for quality control, yet existing methods often struggle with complex backgrounds and shape-specific defects. In this paper, we propose an improved fabric defect detection model based on YOLOv4. To enhance the detection of strip defects, we introduce a Strip Perception Module (SPM) that improves feature capture through multi-scale convolution. We further enhance the spatial pyramid pooling fast (SPPF) by integrating a squeeze-and-excitation mechanism, resulting in the SE-SPPF module, which better integrates spatial and channel information for more effective defect feature extraction. Additionally, we propose a novel focal enhanced complete intersection over union (FECIoU) metric with adaptive weights, addressing scale differences and class imbalance by adjusting the weights of hard-to-detect instances through focal loss. Experimental results demonstrate that our model achieves a 0.8-8.1% improvement in mean average precision (mAP) on the Tianchi dataset and a 1.6-13.2% improvement on our custom dataset, outperforming other state-of-the-art methods.
Index Terms:
fabric defect detection, multi-scale convolution, squeeze-and-excitation networks, deep learning, intersection over union loss function, fabric defect dataset
@@ I Introduction
Traditional fabric defect detection [1, 2, 3] relies heavily on visual inspection by human experts, a process that is time-consuming, labor-intensive, and prone to errors, particularly when defects are small or contrast is low. This method often produces subjective and difficult-to-quantify results, leading to high defect rates and unreliable assessments. As a result, computer vision-based defect detection algorithms [4, 5, 6, 7, 8, 9, 10] have begun to emerge and develop. However, general object detection algorithms struggle with the complex backgrounds of fabric defects and their varied aspect ratios. Thus, adapting to the large-scale variations of fabric defects and distinguishing complex backgrounds are key challenges in improving the performance of fabric defect detection.
Modern fabric defect detection algorithms are generally divided into two categories: two-stage and single-stage methods. The two-stage method, such as Cascade Region-based Convolutional Neural Networks (Faster R-CNN) [11], improves accuracy and speed through cascaded detection. However, it may struggle with detecting multiple defects or misidentifying them. Similarly, the Convolutional Neural Network-based Mobile-Unet method [12] faces similar limitations. Recently, diffusion models have also gained attention in various vision tasks, including fabric defect detection, due to their ability to generate high-quality outputs and handle complex visual patterns. For instance, IMAGPose [13] and IMAGDressing [14] have demonstrated the potential of diffusion models for pose-guided image synthesis and customizable virtual dressing. Additionally, advancements in progressive conditional diffusion models [15] and rich-contextual conditional diffusion models [16] have shown promise in enhancing the consistency and realism of generated images, offering a potential direction for fabric defect detection in more complex scenarios.
@@
The single-stage method, derived from the YOLO [17] framework, has shown promise. For example, the enhanced YOLOv3 [18] model [19] improves detection through an attention mechanism and negative sample weighting but remains insufficient for accurately detecting complex defect types. The YOLOv5 [20] algorithm [21] enhances feature representation by combining adaptive pooling with an attention module and optimizing the loss function. However, its accuracy remains limited in handling specific defect types and complex scenarios.
In response to these challenges, we propose a fabric defect detection model based on the improved YOLOv4 [22] model. While retaining the speed advantages of single-stage models, we introduce a Strip Perception Module (SPM) that incorporates multi-scale convolution to significantly enhance the model’s feature capture and extraction capabilities for strip defects. To improve the ability to distinguish between complex backgrounds and defects, we propose an enhanced Squeeze-and-Excitation Spatial Pyramid Pooling Fast (SE-SPPF), which fully integrates spatial and channel features. Additionally, to address the wide range of target box scales for different defect types, we introduce the Focal Enhanced Complete Intersection over Union (FECIoU) metric. This novel approach dynamically adjusts weights for difficult-to-detect instances, improving the model’s adaptability to target boxes with large aspect ratios.
The key contributions of this paper are as follows:
@@
    A multi-scale convolutional SPM is introduced into the YOLOv4 backbone to improve feature capture and extraction for strip defects.
@@
    We propose FECIoU, an improved version of CIoU, which incorporates a focal weighting mechanism to reduce the impact of scale variations in fabric defects, improving both detection efficiency and accuracy.
@@ II Related Work
II-A Fabric Defect Detection Algorithms
Modern fabric defect detection methods are mainly divided into two categories: two-stage and single-stage approaches. The two-stage methods, such as Faster R-CNN [11], utilize cascade detection to improve accuracy and speed. However, they may struggle with detecting multiple defects in a single fabric sample. The CNN-based Mobile-Unet [12], which replaces U-Net’s encoding block with MobileNetV2, achieves impressive accuracy (99.75% on YID, 98.80% on FID), but still faces limitations in handling various defect types. Single-stage methods, particularly those based on the YOLO framework, have gained popularity. Enhanced YOLOv3 [18] improves fabric defect detection by adding an attention mechanism and negative sample weighting [19]. While effective, it still underperforms in detecting complex defects. YOLOv5 [20] improves feature representation through adaptive pooling and an attention module [21], but faces challenges in complex scenarios. To address these issues, we propose an improved YOLOv4-based model with a Strip Perception Module (SPM) that enhances feature extraction for strip defects, retaining the speed advantage of single-stage detection.
II-B Attention Mechanism
The attention mechanism [23, 24] enhances model performance by focusing on relevant spatial, channel, or hybrid features. Spatial attention methods like SAM [25] and RANet [26] prioritize key regions in the spatial domain, improving the capture of spatial dependencies. RANet uses a relation module to model feature interactions, leveraging attention or graph convolutions. Channel attention, exemplified by SENets [27], introduces a squeeze-and-excitation (SE) block that reweights feature channels to highlight important features. This mechanism improves representational power without significantly increasing computational cost. For fabric defect detection, we propose an enhanced spatial pyramid pooling fast (SE-SPPF) that integrates SENetv2 [28] for better multi-scale feature fusion, addressing the complexity and variation of defect shapes.
II-C Loss Function
@@
In object detection, the loss function quantifies the difference between predicted and ground truth bounding boxes. Intersection over Union (IoU) [29] is commonly used to measure this overlap. The IoU loss encourages the model to align predicted boxes with ground truth. The Generalized IoU (GIoU) [30] extends IoU by addressing scale and offset mismatches, providing more reliable localization, but it can be ineffective for boxes with significant overlap. Distance IoU (DIoU) [31] refines GIoU by incorporating centroid distance, improving localization accuracy. However, DIoU does not account for size variations between objects. Complete IoU (CIoU) [31] incorporates centroid distance, overlap area, and angular difference, making it more effective for rotated boxes. However, for fabric defect detection, where target aspect ratios vary significantly, basic IoU can lead to errors. To address this, we propose an improved version of CIoU (FECIoU), which adjusts for scale differences and enhances detection accuracy for targets with varying aspect ratios.
Figure 2: Strip Perception Module
@@ III-A Overview
This paper presents a fabric defect detection method based on YOLOv11, addressing the challenges of complex defect shapes and the need for high detection accuracy and real-time performance. The proposed method incorporates a strip perception module (SPM) and a squeeze-and-excitation spatial pyramid pooling fast (SE-SPPF). As shown in Fig.1, this approach enhances YOLOv11 by maintaining high detection accuracy while meeting real-time constraints, achieving significant improvements in fabric defect detection.
The SPM leverages strip convolution to extract strip defect features through intensive interactions with convolutions of various shapes, improving the model’s precision in detecting and positioning strip defects. To enhance background discrimination and texture information extraction, the spatial pyramid pooling is re-designed as SE-SPPF, combining the channel attention mechanism of SENetv2. This module optimally utilizes both channel and spatial information to refine background discrimination and defect feature extraction. Additionally, a novel loss function, focal enhanced complete intersection over union (FECIoU), is introduced to address the issue of large-scale variations in target boxes. FECIoU assigns higher weights to samples with lower IoU, ensuring the model focuses on these challenging samples during training, thus improving detection efficiency and accuracy.
@@ III-B Strip Perception Module
In the task of fabric defect detection, the complex shape and large size variation of defect features affect the accuracy of detection. Multi-scale convolution can effectively capture features at different scales in the feature map, especially when facing long strip-shaped defects that occur frequently in fabric operations. Multi-scale convolution can more effectively extract defect features. The specific design is shown in Fig.2.
This paper proposes SPM. First, two convolution blocks of 1×1 and 3×3 are used to minimize the number of channels, and then multi-scale (1×3, 3×1, 3×3) convolution operations are performed using branch parallelism. The resulting feature maps are densely stacked using concat, and then a 1x1 convolution kernel is used to extract important features from the convolutions of different scales. Finally, a residual structure was introduced to improve the stability and effectiveness of training. While maintaining the depth of the network, information transmission and gradient flow are ensured. In summary, SPM can effectively extract the features of strip defects and improve the accuracy of the model.
@@ III-C Squeeze and Excitation Spatial Pyramid Pooling Fast
Fabric defects usually exhibit multiple features. In order to eliminate some noise, make the features more robust, and help the model better capture the overall structure and texture of the image, SE-SPPF introduces SENetv2 to more reasonably assign weights to each channel. Combined with the multi-scale fusion in SPPF space, it strengthens the model’s ability to extract features from both spatial and channel perspectives. The specific design is shown in Fig.3. This paper proposes SE-SPPF. First, the feature map is weighted by SENetv2 to the channel, and then the channel number is adjusted using a 1x1 convolution and input to SPPF. The four feature maps of different scales obtained by SPPF are concatenated using a residual structure and the weighted feature map Concat after feature extraction using a 1x1 convolution. Finally, features are further extracted using two convolutions of 1x1 and 3x3.
@@ III-D Focal Enhanced Complete Intersection over Union
The span of the defect detection box for different types of fabric defects is very large, especially for defects that appear in the form of stripes, which are several times or even more than the length and width of most target detection objects. Therefore, this paper proposes FECIoU, which uses a focal weight mechanism to make the model pay more attention to difficult-to-detect objects during training. Equation 1 is the formula for FECIoU, where $(1-IoU)^{\gamma}$is the weight value for CIoU and $\gamma$ is a manually set parameter. In Equation 2 ,$\rho^{2}(b,b^{g})$is the squared Euclidean distance between the centers of the predicted and ground truth boxes, calculated as shown in Equation 3, and $c$ is the diagonal length of the minimum bounding box. $\alpha v$ is a penalty term for the aspect ratio difference, and the specific calculation method is shown in Equations 4 and 5 . $w^{g},h^{g},w,$ and $h$ are the width and height of the predicted frame and the actual frame, respectively.
@@
Figure 3: Squeeze-and-Excitation Spatial Pyramid Pooling Fast
@@ IV-A Datasets
Tianchi fabric dataset Tianchi fabric dataset [32], provided by Alibaba’s Tianchi platform, is a significant resource for fabric defect detection research. It comprises high-resolution fabric images with detailed annotations of various defect types, such as holes, stains, wrinkles, color shades, and missing threads. The dataset, consisting of thousands to tens of thousands of images, is designed to facilitate the development and validation of defect detection algorithms and automated quality inspection systems in the fabric industry.

### v3

@@ Abstract
Defect detection in fabrics is critical for quality control, yet existing methods often struggle with complex backgrounds and shape-specific defects. In this paper we propose SPFFNet, an improved fabric defect detection model based on the YOLOv11 framework. To enhance the detection of strip defects, we introduce a Strip Perception Module (SPM) that improves feature capture through multi-scale convolution. We further enhance the spatial pyramid pooling fast (SPPF) by integrating a squeeze-and-excitation mechanism, resulting in the SE-SPPF module, which better integrates spatial and channel information for more effective defect feature extraction. Additionally, we propose a novel focal enhanced complete intersection over union (FECIoU) metric with adaptive weights, addressing scale differences and class imbalance by adjusting the weights of hard-to-detect instances through focal loss. Experimental results demonstrate that our model achieves a 0.8-8.1% improvement in mean average precision (mAP) on the Tianchi dataset and a 1.6-13.2% improvement on our custom dataset, outperforming other state-of-the-art methods.
Figure 1: Network structure of the proposed method
@@ I Introduction
Traditional fabric defect detection [weng_2024_enhancing, li_2024_lrfpn, qiao_2022_a] relies heavily on visual inspection by human experts, a process that is time-consuming, labor-intensive, and prone to errors, particularly when defects are small or contrast is low. This method often produces subjective and difficult-to-quantify results, leading to high defect rates and unreliable assessments. As a result, computer vision-based defect detection algorithms, have begun to emerge and develop. However, general object detection algorithms struggle with the complex backgrounds of fabric defects and their varied aspect ratios. Thus, adapting to the large-scale variations of fabric defects and distinguishing complex backgrounds are key challenges in improving the performance of fabric defect detection.
Modern fabric defect detection algorithms are generally divided into two categories: two-stage and single-stage methods. The two-stage method, such as Zhao et al.  [jia_2022_fabric] proposed a transfer learning-based Faster Region-based Convolutional Neural Network (Faster R-CNN), enhancing fabric defect detection accuracy via a cascaded module. However, it faces challenges in training efficiency, computational cost, and generalizability to complex textures and diverse defects.
@@
The single-stage method, derived from the YOLO [redmon_2016_you] framework, has shown promise. For example, the enhanced YOLOv3 [redmon_2018_yolov3] model [jing_2020_fabric] improves detection through an attention mechanism and negative sample weighting but remains insufficient for accurately detecting complex defect types. The YOLOv5 [jocher_2020_ultralyticsyolov5] algorithm [liu_2023_an] enhances feature representation by combining adaptive pooling with an attention module [zhao2024athlete, luo2025tennis] and optimizing the loss function. However, its accuracy remains limited in handling specific defect types and complex scenarios.
In response to the aforementioned challenges, we propose SPFFNet, a novel architecture built upon the YOLOv11 framework [jocher_2023_yolov8]. While preserving the inference efficiency characteristic of single-stage detectors, SPFFNet introduces a Strip Perception Module (SPM) that leverages multi-scale convolution to substantially enhance the network’s capability for fine-grained feature extraction and representation of strip defects. To further improve discrimination between complex background textures and subtle defect regions, we design an enhanced Squeeze-and-Excitation Spatial Pyramid Pooling Fast (SE-SPPF) module, which effectively integrates spatial and channel-wise information to achieve more comprehensive contextual understanding. Moreover, to address the substantial variation in bounding-box scales across different defect categories, we propose a Focal Enhanced Complete Intersection over Union (FECIoU) metric. This metric dynamically reweights difficult-to-detect samples, thereby improving robustness and adaptability to targets with extreme aspect ratios.The main contributions of this work are summarized as follows:
@@
    A multi-scale convolutional SPM is introduced into the YOLOv11 backbone to improve feature capture and extraction for strip defects.
@@
    We propose FECIoU, which incorporates a focal weighting mechanism to reduce the impact of scale variations in fabric defects.
@@ II Related Work
II-A Object Detection
Object detection has long been an active research area, with numerous methods developed to enhance detection accuracy and efficiency. Early approaches such as R-CNN [girshick_2014_rich] combined region proposals with CNN-based feature extraction, while Fast R-CNN [girshick_2015_fast] and Faster R-CNN [ren_2017_faster] improved both speed and accuracy through shared convolutional features and the introduction of the Region Proposal Network (RPN). Subsequently, one-stage detectors like YOLO [redmon_2016_you] and SSD [liu_2016_ssd] achieved real-time object detection by formulating detection as a single regression problem. Despite these advances, challenges remain in detecting small objects and handling diverse object scales and shapes.
@@ II-A Object Detection
In object detection, the loss function quantifies the difference between predicted and ground truth bounding boxes. Intersection over Union (IoU) [jiang_2018_acquisition] is commonly used to measure this overlap. The IoU loss encourages the model to align predicted boxes with ground truth. The Generalized IoU (GIoU) [rezatofighi_2019_generalized] extends IoU by addressing scale and offset mismatches, providing more reliable localization, but it can be ineffective for boxes with significant overlap. Distance IoU (DIoU) [zheng_2019_distanceiou] refines GIoU by incorporating centroid distance, improving localization accuracy. However, DIoU does not account for size variations between objects. Complete IoU (CIoU) [zheng_2019_distanceiou] incorporates centroid distance, overlap area, and angular difference, making it more effective for rotated boxes. However, for fabric defect detection, where target aspect ratios vary significantly, basic IoU can lead to errors. To address this, we propose an improved version of CIoU (FECIoU), which adjusts for scale differences and enhances detection accuracy for targets with varying aspect ratios.
II-B Fabric Defect Detection Algorithms
Modern fabric defect detection methods are mainly divided into two categories: two-stage and single-stage approaches. The two-stage methods, such as Zhao et al.  [jia_2022_fabric] proposed method, which integrates transfer learning and an improved Faster R-CNN with Residual Network with 50 layers (ResNet50), Feature Pyramid Network (FPN), Region of Interest Align (ROI Align), significantly enhances detection accuracy and robustness for fabric defect detection by employing a cascaded module to refine localization precision. Single-stage methods, particularly those based on the YOLO framework, have gained popularity. Enhanced YOLOv3 [redmon_2018_yolov3] improves fabric defect detection by adding an attention mechanism and negative sample weighting [jing_2020_fabric]. While effective, it still underperforms in detecting complex defects. YOLOv5 [jocher_2020_ultralyticsyolov5] improves feature representation through adaptive pooling and an attention module [liu_2023_an], but faces challenges in complex scenarios. To address these issues, we propose an improved YOLOv4-based model with a Strip Perception Module (SPM) that enhances feature extraction for strip defects, retaining the speed advantage of single-stage detection.
II-C Attention Mechanism
The attention mechanism [shen_2023_pedestrianspecific, shen_2023_triplet] enhances model performance by focusing on relevant spatial, channel, or hybrid features. Spatial attention methods like SAM [zhu_2019_an] and RANet [shao_2023_ranet] prioritize key regions in the spatial domain, improving the capture of spatial dependencies. RANet uses a relation module to model feature interactions, leveraging attention or graph convolutions. Channel attention, exemplified by SENets [hu_2018_squeezeandexcitation], introduces a squeeze-and-excitation (SE) block that reweights feature channels to highlight important features. This mechanism improves representational power without significantly increasing computational cost. For fabric defect detection, we propose an enhanced spatial pyramid pooling fast (SE-SPPF) that integrates SENetv2 [narayanan_2023_senetv2] for better multi-scale feature fusion, addressing the complexity and variation of defect shapes.
Figure 2: (a) illustrates the overall architecture of the Strip Perception Module, while (b) presentsdetailed fusion operations within the Dense Connections component. The Dense Connectionsfacilitate comprehensive integration of diverse strip-wise features.
@@ III-A Overview
This paper presents a fabric defect detection method based on YOLOv11, addressing the challenges of complex defect shapes and the need for high detection accuracy and real-time performance. The proposed method incorporates a strip perception module (SPM) and a squeeze-and-excitation spatial pyramid pooling fast (SE-SPPF). As shown in Fig.1, this approach enhances YOLOv11 by maintaining high detection accuracy while meeting real-time constraints, achieving significant improvements in fabric defect detection. The SPM leverages strip convolution to extract strip defect features through intensive interactions with convolutions of various shapes, improving the model’s precision in detecting and positioning strip defects. To enhance background discrimination and texture information extraction, the spatial pyramid pooling is re-designed as SE-SPPF, combining the channel attention mechanism of SENetv2. This module optimally utilizes both channel and spatial information to refine background discrimination and defect feature extraction. Additionally, a novel loss function, focal enhanced complete intersection over union (FECIoU), is introduced to address the issue of large-scale variations in target boxes. FECIoU assigns higher weights to samples with lower IoU, ensuring the model focuses on these challenging samples during training, thus improving detection efficiency and accuracy.
@@ III-B Strip Perception Module
In the task of fabric defect detection, the complex shape and large size variation of defect features affect the accuracy of detection. Multi-scale convolution can effectively capture features at different scales in the feature map, especially when facing long strip-shaped defects that occur frequently in fabric operations. Multi-scale convolution can more effectively extract defect features. The specific design is shown in Fig.2. This paper proposes SPM. First, two convolution blocks of 1x1 and 3x3 are used to minimize the number of channels, and then multi-scale (1x3, 3x1, 3x3) convolution operations are performed using branch parallelism. The resulting feature maps are densely stacked using concat, and then a 1x1 convolution kernel is used to extract important features from the convolutions of different scales. Finally, a residual structure was introduced to improve the stability and effectiveness of training. While maintaining the depth of the network, information transmission and gradient flow are ensured. In summary, SPM can effectively extract the features of strip defects and improve the accuracy of the model.
Figure 3: (a) illustrates the overall architecture of the Squeeze-and-Excitation Spatial Pyramid Pooling Fast (SE-SPPF), while (b) provides a detailed breakdown of the SENetV2 module’s processing pipeline. The SENetV2 module effectively harmonizes multi-scale features to extract the most discriminative characteristics for defect detection.
mAP@0.5/% Method Knot Triple Wire Coarse Pick Broken Spandex Warp Knot Weft Shrink Hole Stain All GFLOPs Params YOLOv5 [jocher_2020_ultralyticsyolov5] 61.6 77.2 59.4 76.8 45.8 46.9 83.7 43.5 61.9 5.8 2183224 YOLOv6 [li_2022_yolov6] 58.1 78 52.8 68.1 47.1 30.8 82.3 44.5 57.7 11.5 4155816 YOLOv8 [jocher_2023_yolov8] 65.9 78.8 60.5 76.3 51.3 40.1 81.6 59.9 64.3 6.8 2685928 YOLOv9t [wang_2024_yolov9] 65.4 80.4 59.8 71.8 52.6 46.9 83.3 62.9 65.4 6.4 1731384 YOLOv9s [wang_2024_yolov9] 66.0 82.0 54.3 76.6 54.4 46.7 79.7 64.4 65.5 22.1 6196744 YOLOv10n [wang_2024_yolov10] 59.3 77.4 57.7 69.4 41.5 39.2 81.7 57.7 60.5 8.2 2697536 YOLOv11n [jocher_2023_yolov8] 64.4 80.0 64.3 76.1 48.1 43.7 80.5 62.9 65.0 6.3 2583712 SPFFNet (Ours) 64.5 80.5 63.5 74.6 49.0 43.9 83.7 66.4 65.8 6.8 2858951
@@ III-C Squeeze and Excitation Spatial Pyramid Pooling Fast
Fabric defects usually exhibit multiple features. In order to eliminate some noise, make the features more robust, and help the model better capture the overall structure and texture of the image, SE-SPPF introduces SENetv2 to more reasonably assign weights to each channel. Combined with the multi-scale fusion in SPPF space, it strengthens the model’s ability to extract features from both spatial and channel perspectives. The specific design is shown in Fig.3. This paper proposes SE-SPPF. First, the feature map is weighted by SENetv2 to the channel, and then the channel number is adjusted using a $1\times 1$ convolution and input to SPPF. The four feature maps of different scales obtained by SPPF are concatenated using a residual structure and the weighted feature map Concat after feature extraction using a $1\times 1$ convolution. Finally, features are further extracted using two convolutions of $1\times 1$ and $3\times 3$.
@@ III-D Focal Enhanced Complete Intersection over Union
The span of the defect detection box for different types of fabric defects is very large, especially for defects that appear in the form of stripes, which are several times or even more than the length and width of most target detection objects. Therefore, this paper proposes FECIoU, which uses a focal weight mechanism to make the model pay more attention to difficult-to-detect objects during training. Equation 1 is the formula for FECIoU, where $(1-IoU)^{\gamma}$is the weight value for CIoU and $\gamma$ is a manually set parameter. In Equation 2 $\rho^{2}(b,b^{g})$is the squared Euclidean distance between the centers of the predicted and ground truth boxes, calculated as shown in Equation 3, and $c$ is the diagonal length of the minimum bounding box. $\alpha v$ is a penalty term for the aspect ratio difference, and the specific calculation method is shown in Equations 4 and 5 . $w^{g},h^{g},w,$ and $h$ are the width and height of the predicted frame and the actual frame, respectively.
@@ IV-A1 Tianchi fabric dataset
Tianchi fabric dataset [tianchi_2020_smart], provided by Alibaba’s Tianchi platform, is a significant resource for fabric defect detection research. It comprises high-resolution fabric images with detailed annotations of various defect types, such as holes, stains, wrinkles, color shades, and missing threads. The dataset, consisting of thousands to tens of thousands of images, is designed to facilitate the development and validation of defect detection algorithms and automated quality inspection systems in the fabric industry.
TABLE II: Comparison of the performance of the proposed improved model with multiple SOTA on the produced dataset
mAP@0.5/% Method Missing Stitches Broken Holes Stain Broken Seam Broken Stitches All GFLOPs Params YOLOv5 [jocher_2020_ultralyticsyolov5] 85.4 73.4 99.5 80.2 75.9 82.9 5.8 2182639 YOLOv6 [li_2022_yolov6] 83.0 68.9 99.5 80.2 55.5 77.4 11.5 4155519 YOLOv8 [jocher_2023_yolov8] 93.9 78.2 99.5 82.0 88.1 88.3 6.8 2685343 YOLOv9t [wang_2024_yolov9] 89.1 76.3 99.5 82.1 85.8 86.5 6.4 1730799 YOLOv9s [wang_2024_yolov9] 91.7 80.2 99.5 81.2 91.8 88.9 22.1 6195583 YOLOv10n [wang_2024_yolov10] 89.5 76.8 99.5 78.6 85.9 86.1 8.3 2696336 YOLOv11n [jocher_2023_yolov8] 93.1 79.4 99.5 83.8 89.3 89.0 6.3 2583127 SPFFNet (Ours) 95.3 83.5 99.5 81.1 93.5 90.6 6.8 2858951
IV-A2 Produced dataset